OpenStudy (anonymous):

What is X+5y=5 in slope intercept form?

OpenStudy (anonymous):

Well you would subtract x from both sides and then you would have 5y=x-5, Then you would divide both sides by 5 and then you would have y=1/5X-1

OpenStudy (anonymous):

Thank you! This is a bit confusing to me and I have a paper of these due next period. I've got 5 to go. Ugh

OpenStudy (anonymous):

If u want I can keep helping I love math

OpenStudy (anonymous):

I'm terrible at it. I still have 16.) -2x+5y= -10 17.) 3x-2y=-12 18.) 2x-3y = 0 19.) 2x+y=5 20.) 4x-5y=-25 21.) 2x-4y=12

OpenStudy (anonymous):

16)y=2/5X-2

OpenStudy (anonymous):

Thank you so much for the help

OpenStudy (anonymous):

17)y=3/2X-6 18)y=2/3X 19)y=-2x+5 20)y=4/5X+5 21)y=1/2X-3 I hope I helped a medal will be appreciated

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem, thank you so much

OpenStudy (anonymous):

Yup I love math